OpenStudy (anonymous):

Find the surface area of an orange with a 3'' radius.

OpenStudy (aravindg):

you can take the orange as a sphere .Then do you know the formula for surface area of a sphere?

OpenStudy (ryan123345):

\[4 \pi r^2\]

OpenStudy (ryan123345):

where r is the radius :)

OpenStudy (anonymous):

Yea you can take it as a sphere. SA=4pirsqrd

OpenStudy (aravindg):

here r=3 substitute !!

OpenStudy (anonymous):

I know but I find it some what confusing..

OpenStudy (aravindg):

r=3 /r^2= ?

OpenStudy (anonymous):

No like do I multiply pi at the end?

OpenStudy (aravindg):

yep :)

OpenStudy (anonymous):

So instead of 28 pi..

OpenStudy (aravindg):



OpenStudy (anonymous):

87.92?

OpenStudy (aravindg):

nope

OpenStudy (aravindg):

9 X 4=36

OpenStudy (anonymous):

WOW my common multiplication instincts are on the floor...

OpenStudy (anonymous):

This is embarrassing.. LOL

OpenStudy (aravindg):

lol it happens sometimes !

OpenStudy (anonymous):